OpenStudy (anonymous):

Solve 1/7 + 5/6. Then use estimation to check for reasonableness. Show your work.

OpenStudy (anonymous):

6/7*6 + 5*7/7*6 = 41/42 An approximation is that 1/7 is a bit lower than 1/6 and 1/6+5/6 = 1 41/42 is a bit lower than 1.